Enter the quadratic equation in standard form. x^2 - 10 + 3x = (x - 2)^2 + 1

Answer by rapaljer(4671)

x^2 -10+3x=x^2 -4x+4 + 1

Subtract x^2 from each side, and all the x^2 terms are gone. Therefore, this is NOT a quadratic equation after all.

-10+3x= -4x+5

Add +4x to each side:
-10+3x+4x=-4x+4x+5
-10+7x=5

Add +10 to each side:
7x= 15

Divide both sides by 7:
x=15/7

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus